For the Fourier series in Problem 1, deduce a

series for

π

4

at the point where x =

π

2


1
Expert's answer
2021-08-11T12:53:47-0400

For the Fourier series in Problem 1, deduce a series for "\dfrac{\pi}4" at the point where "x= \dfrac{\pi}2"


"f(x)= \dfrac{a_0}2+ \Sigma a_n \sin \dfrac{nπx}{2}+ \Sigma b_n\cos\dfrac{nπx}{2}"

but bn = 0

Hence, "f(x)= \dfrac{a_0}2+ \Sigma a_n \sin \dfrac{nπx}{2}"

When "\dfrac{\pi}2", f(x) = 2, hence;


"2 = \dfrac8{\pi}\sin\dfrac{\pi}2+\dfrac8{3\pi}\sin\dfrac{3\pi}2+ \dfrac8{5\pi}\sin\dfrac{5\pi}2+..."


i.e

"2 =\dfrac8{\pi}(1-\dfrac13+\dfrac15-\dfrac17+...)"


"\dfracπ4 = 1-\dfrac13+\dfrac15-\dfrac17+..."